# Closed mean curvature flows with asymptotically conical singularities

### v2

@@ Abstract.
In this paper, we prove that for any asymptotically conical self-shrinker, there exists an embedded closed hypersurface such that the mean curvature flow starting from it develops a singularity modeled on the given shrinker. The main technique is the Ważewski box argument, used by Stolarski in the proof of the corresponding theorem in the Ricci flow case. As a corollary, our construction, combined with the works of Angenent–Ilmanen–Velázquez and Chodosh–Daniels-Holgate–Schulze, implies the existence of fattening level set flows starting from smooth embedded closed hypersurfaces. This addresses a question posed by Evans–Spruck and De Giorgi.
@@ 1. Introduction
In [W02], White announced a result joint with Ilmanen about the existence of a compact smooth embedded surface in $\mathbb{R}^{3}$ for which the uniqueness of enhanced varifold solutions of mean curvature evolution fails. He described the proof ideas in [W02, Section 4], which are completely different from our approach to Corollary 1.1. Combining their result with Corollary 1.1, we now have $n$-dimensional embedded closed examples for $n\in\left\{2,3,4,5,6\right\}.$
@@
The proof of Theorem 1 follows the outline of the argument of the corresponding construction in the Ricci flow case by Stolarski [Sto]. More precisely, we used the Ważewski box argument to find a good perturbation using the eigenfunctions of the stability operator of the Gaussian functional. The goal is to make the mean curvature flow starting from a perturbation of a doubling of an asymptotically conical shrinker develop a singularity modeled on the chosen shrinker itself.
@@
There are three main difficulties in the case of mean curvature flows. The first one is that we have to write the flow as a normal graph over the shrinker to use the tools from analysis, while in the Ricci flow case [Sto], one only needs to pull back tensors. The second difficulty is that we have to guarantee the embeddedness of the modified doubling of an asymptotically conical shrinker and control the curvature bound on the gluing part, while in the Ricci flow case, we can use cutoff functions to modify the metric directly. The third difficulty is that the shrinker quantity is more nonlinear than the Ricci curvature and hence complicates the arguments.
@@
This paper is organized as follows. In Section 2, we provide preliminary notations and results for asymptotically conical self-shrinkers and talk about the evolution equations of related immersions. In Section 3, we give an explicit construction of an embedded doubling of an asymptotically conical shrinker. In Sections 4 and 5, we talk about the interior estimates and constructions of barrier functions to get the preservation and the improvement of $C^{2}$ bounds. In Section 6, we use the Ważewski box argument to finish the proof of Theorem 1 and prove Corollary 1.1.
@@ Acknowledgement
The authors are grateful to Prof. Toby Colding and Prof. Bill Minicozzi for their constant support and several inspiring conversations. They would like to thank Letian Chen, Otis Chodosh, Alec Payne, Max Stolarski, and Ao Sun for valuable discussions, and to Alec Payne for bringing the work of Angenent, Ilmanen, and Velázquez to their attention and suggesting the application regarding fattening. During the project, Lee was partially supported by NSF Grant DMS 2005345 and Zhao was supported by NSF Grant DMS 1812142 and NSF Grant DMS 1811267.
@@ 2. Preliminaries facts and evolution equations
2.1. Stability operators on asymptotically conical shrinkers
@@
There exists an $L^{2}_{W}$-orthonormal basis $\{h_{i}\}_{i\in\mathbb{N}}$ of $L^{2}_{W}(\Sigma)$ such that for all $i\in\mathbb{N},$ $h_{i}$ is an eigenfunction of $\mathcal{L}$ with eigenvalue $\lambda_{i},$ that is, $\mathcal{L}h_{i}=-\lambda_{i}h_{i}.$ The eigenvalues $\lambda_{i}$’s satisfy $\lambda_{1}<\lambda_{2}\leq\lambda_{3}\leq\cdots$ tending to +,$\infty$,+ ∞ , and each eigenspace is finite dimensional.
@@
$|h(x)|\leq C|x|^{\lambda+\delta+\frac{1}{2}}.$
@@ 2.2. Evolution equation of immersions
$(\partial_{t}F_{0})(\varphi_{t}(x),t)+dF_{0}(\partial_{t}\varphi_{t}(x))(\varphi_{t}(x),t)$
$=$ $\sqrt{1-t}\,\partial_{t}{h}(x,t)\textbf{n}_{F}(x)-\frac{1}{2\sqrt{1-t}}\left(F(x)+{h}(x,t)\textbf{n}_{F}(x)\right).$
@@
$(\partial_{t}F_{0})(\varphi_{t}(x),t)+dF_{0}(\partial_{t}\varphi_{t}(x))(\varphi_{t}(x),t)+\frac{1}{2\sqrt{1-t}}F(x)$
$=$ $(\sqrt{1-t}\,\partial_{t}{h}(x,t)-\frac{1}{2\sqrt{1-t}}{h}(x,t))\textbf{n}_{F}(x).$
As we assume that locally $\frac{1}{\sqrt{1-t}}F_{1}$ can be written as a graph over $F$, we have
$\langle\textbf{n}_{F_{0}}(\varphi_{t}(x)),\textbf{n}_{F}(x)\rangle\neq 0.$
We denote
@@
(2.7) $|dF_{0}(\partial_{t}\varphi_{t}(x))|\leq\frac{\left|(\partial_{t}F_{0})(\varphi_{t}(x),t)+\frac{1}{2\sqrt{1-t}}F(x)\right|}{|\langle\textbf{n}_{F}(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle|}.$
@@
Setting $e^{\tau}=\frac{1}{1-t},$ we get
$\partial_{t}=\frac{\partial\tau}{\partial t}\partial_{\tau}=\frac{1}{1-t}\partial_{\tau},$
and thus,
@@
$\hat{h}_{ab}$ $=\langle\partial_{a}F,\partial_{b}(h\,\textbf{n}_{F})\rangle+\langle\partial_{a}(h\,\textbf{n}_{F}),\partial_{b}F\rangle+\langle\partial_{a}(h\,\textbf{n}_{F}),\partial_{b}(h\,\textbf{n}_{F})\rangle=2hA_{ab}+\langle\partial_{a}h,\partial_{b}h\rangle+h^{2}A_{ia}A_{bj}g^{ij},$
@@
$\|C_{1}(F,\nabla F,\nabla^{2}F,h,\nabla h)\|_{C^{0}}+\sum_{a,b}\|C_{2,ab}(F,\nabla F,\nabla^{2}F,h,\nabla h)\|_{C^{0}}+\|C_{3,ab}(F,\nabla F,\nabla^{2}F,h,\nabla h)\|_{C^{0}}\leq C_{0}\|h\|_{C^{1}}.$
Also, note that $\Delta^{g}\textbf{n}_{F}=-|A|^{2}\textbf{n}_{F}+\nabla^{g}H$ and $A_{ij,k}=A_{ik,j}$ in $\mathbb{R}^{n+1}$. Here, we use the notation $\textbf{H}=\Delta^{g}F$ and $H=-\textbf{H},\textbf{n}_{F}\rangle=g^{ij}A_{ij},$ so in normal coordinates, we have
@@ 3. Constructions of doubling of asymptotically conical shrinkers
As in (2.1), denoting the asymptotic cone of $\Sigma$ by $\mathcal{C}$, from [W14, CS], we know that outside a large compact set, we can parameterize $\Sigma\setminus B_{R}$ by
@@
over $\mathcal{C}\setminus B_{R}$, where $\theta$ is the parametrization for the link of $C$ and n is the normal vector of $\mathcal{C}$ at $(r,\theta)$.
@@
$F_{3}(r,\theta)=\begin{cases}r\theta+\eta(\frac{r-R}{R})f(r,\theta)\textbf{n}&\text{if }r\leq 2R\\ r\cdot\frac{\left(\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)+\frac{1}{3}\right)\theta+\left(\frac{2}{3}-\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)\right)\mathbf{v}}{\left|\left(\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)+\frac{1}{3}\right)\theta+\left(\frac{2}{3}-\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)\right)\mathbf{v}\right|}&\text{if }r\geq 2R\end{cases}.$
@@
From the constructions, the $F_{3}$ part lies in a hemisphere and controls the distance to origin. Thus, we know that $F$ is globally an embedding, and after reflecting it with respect to the hyperplane $\{x\in\mathbb{R}^{n+1}\,|\,\langle x,\mathbf{v}\rangle=B+R\},$ we get a closed embedded hypersurface.
@@ 4. Interior estimates
$\|h\|_{L^{\infty}(\Sigma\times[0,\tau_{1}])}$ $<\varepsilon,$
$\|h(\tau)\|_{L_{\frac{|f|^{2}}{4}}^{2}(\Sigma)}$ $\leq\mu e^{\lambda_{*}(\tau+\tau_{0})}\,\,\text{ for all}\,\,\tau\in[0,\tau_{1}],\text{ and}$
$\sum_{l=0}^{4n+3}\left|\nabla^{l}h\right|_{g}(x,0)$ $\leq C_{0}\bar{p}e^{\lambda_{*}\tau_{0}}\,\,\text{ for all}\,\,x\in\left\{\frac{|F|^{2}}{4}<2\Gamma e^{\tau_{0}}\right\}\text{ and }\tau=0.$
Then for $\Gamma>\Gamma_{0}(n,\Sigma,F)$, $\varepsilon<\varepsilon_{0}(n,\Sigma,F,\Gamma),$ and $\tau_{0}>C(n,\Sigma,F,\lambda_{*},\Gamma,C_{0}\bar{p})$, there exists $C_{1}=C_{1}(n,\Sigma,F,\lambda_{*},\Gamma)<\infty$ such that
$\sum_{l=0}^{2}|\nabla^{g}h|_{g}(x,\tau)\leq C_{1}(\mu+C_{0}\bar{p})e^{\lambda_{*}(\tau+\tau_{0})}\,\,\text{ for all}\,\,(x,\tau)\in\left\{\frac{|F|^{2}}{4}<\Gamma e^{\tau_{0}}\right\}\times[0,\tau_{1}].$
@@
This helps us to get rid of the unbounded coefficient term $-\frac{1}{2}\langle F,\nabla^{g}f\rangle$ in $L$. Also, note that $0\leq\left|\nabla^{g}(\frac{|F|^{2}}{4})\right|^{2}\leq\frac{|F|^{2}}{4}$ gives us that for $s_{1}<s_{2},$
@@
Thus, Proposition 4 implies the large scale interior estimates.
@@
$|h|_{g}$ $\leq C_{0}e^{-k(\tau+\tau_{0})}\left(\frac{|F|^{2}}{4}(x)\right)^{k}\,\,\text{on}\,\,\left\{(x,\tau)\in\Sigma\times(0,\tau_{1}):\,\Gamma e^{\tau_{0}}<\frac{|F|^{2}}{4}<\gamma e^{\tau+\tau_{0}}\right\},\text{ and}$
$\sum_{l=0}^{3}|\nabla^{l}h|_{g}(x,\tau_{0})$ $\leq C_{0}e^{-k\tau_{0}}\left(\frac{|F|^{2}}{4}(x)\right)^{k}\,\,\text{for all}\,\,x\in\left\{\Gamma e^{\tau_{0}}<\frac{|F|^{2}}{4}<\gamma e^{\tau+\tau_{0}}\right\}.$
@@
$\sum_{l=0}^{2}|\nabla^{g}h|_{g}(x,\tau)\leq C_{1}C_{0}e^{-k(\tau+\tau_{0})}\left(\frac{|F|^{2}}{4}(x)\right)^{k}\,\,\text{on}\,\,\left\{(x,\tau)\in\Sigma\times[0,\tau_{1}]:\,4\Gamma e^{\tau_{0}}\leq\frac{|F|^{2}}{4}(x)\leq\frac{1}{4}\gamma e^{\tau+\tau_{0}}\right\}.$
Similarly, on the outer region, we have the following estimate.
@@
$|h|_{g}(x,\tau)$ $\leq\delta\,\,\text{on}\,\,\left\{(x,t)\in\Sigma\times[0,\tau_{1}]:\,\gamma e^{\tau+\tau_{0}}<\frac{|F|^{2}}{4}(x)\right\},$
$\sum_{l=0}^{3}|\nabla^{l}h|_{g}(x,0)$ $\leq\delta\,\,\text{on}\,\,x\in\left\{\gamma e^{\tau_{0}}\leq\frac{|F|^{2}}{4}\right\},\text{ and}$
@@
$\sum_{l=0}^{2}|\nabla^{g}h|_{g}(x,\tau)\leq C\delta\,\,\text{on}\,\,\left\{(x,\tau)\in\Sigma\times[0,\tau_{1}]:\,4\gamma e^{\tau+\tau_{0}}<\frac{|F|^{2}}{4}(x)\right\}.$

### v3

@@ Abstract.
In this paper, we prove that for any asymptotically conical self-shrinker, there exists an embedded closed hypersurface such that the mean curvature flow starting from it develops a singularity modeled on the given shrinker. The main technique is the Ważewski box argument, used by Stolarski in the proof of the corresponding theorem in the Ricci flow case. As a corollary, our construction, combined with the works of Angenent–Ilmanen–Velázquez and Chodosh–Daniels-Holgate–Schulze, implies the existence of fattening level set flows starting from smooth embedded closed hypersurfaces. These provide examples related to a question asked by Evans–Spruck.
@@ 1. Introduction
In [W02], White announced a result joint with Ilmanen about the existence of a compact smooth embedded surface in $\mathbb{R}^{3}$ for which the uniqueness of enhanced varifold solutions of mean curvature evolution fails. He described the proof ideas in [W02, Section 4], which are completely different from our approach to Corollary 1.1. In a recent work [IW], Ilmanen–White improved the result and showed that for any large enough $g,$ there exists a smooth embedded closed surface in $\mathbb{R}^{3}$ of genus $g$ such that the level set flow starting from it develops an asymptotically conical singularity (with two ends) and fattens after that singular time. Their work was based on the resolution of the multiplicity one conjecture by Bamler–Kleiner [BK23]. Combining their results with Corollary 1.1, we now have $n$-dimensional embedded closed fattening examples for $n\in\left\{2,3,4,5,6\right\}.$
There was another recent work by Ketover [K24], where he used min-max methods to construct a genus $g$ asymptotically conical shrinker with two ends and an entropy estimate. Furthermore, the level set flow starting from the asymptotic cone of the shrinker fattens when $g$ is large. As pointed out in [IW], combining Ketover’s result with Theorem 1 and arguing as in the proof of Corollary 1.1 leads to other examples of two-dimensional fattening level set flows starting from smooth embedded closed surfaces.
@@
We briefly mention the strategy to prove Theorem 1. The proof of Theorem 1 follows the outline of the argument of the corresponding construction in the Ricci flow case by Stolarski [Sto]. More precisely, we used the Ważewski box argument to find a good perturbation using the eigenfunctions of the stability operator of the Gaussian functional. The goal is to make the mean curvature flow starting from a perturbation of a doubling of an asymptotically conical shrinker develop a singularity modeled on the chosen shrinker itself. To be more specific, we perturb the doubling by eigenfunctions and consider the mean curvature flow starting from the perturbations. By considering the first time when the flow gets outside of the ”boxes with good estimates”, we can see that the only possibility is that its unstable component hits the boundary. By topological arguments, we can prove that there exists a perturbation which always stays in the box, which means that it develops the corresponding singularity.
An obvious difference between our work and [Sto] is the extrinsic nature of mean curvature flows. When working with the Ricci flow, Stolarski perturbed the metric tensor to achieve the goal of constructing the desired closed Riemannian manifold. In the case of mean curvature flows, we perturb a doubling of a shrinker by, as expected, looking at normal graphs on it. Thus, the role of the perturbing function $h$ in our work is not the role of the perturbing tensor in [Sto]. Instead, since metric tensors come from pairings of first derivatives, perturbing tensors should correspond to the gradient of our perturbing functions. This requires additional work to relate the usual $C^{k}$ norms with the homogeneous norms on an asymptotically conical shrinker.
@@
Other than a slight difference mentioned above, there are three main difficulties in the case of mean curvature flows. The first one is that we have to write the flow as a normal graph over the shrinker to use the tools from analysis, while in the Ricci flow case [Sto], one only needs to pull back tensors. Moreover, we have to guarantee the embeddedness of the modified doubling of an asymptotically conical shrinker and control the curvature bound on the gluing part, while in the Ricci flow case, we can use cutoff functions to modify the metric directly. Finally, the shrinker quantity is more nonlinear than the Ricci curvature and hence complicates the arguments.
We remark that there are constructions of mean curvature flows with singularities modeled on minimal cones, which form another important class of singularities. In [V94], Velázquez constructed mean curvature flows with singularities modeled on the Simons cones. Following his work, more precise information has been studied for flows with such singularities. See, for example, [GS18, S23V, Liu].
@@
This paper is organized as follows. In Section 2, we provide preliminary notations and results for asymptotically conical self-shrinkers and talk about the evolution equations of related immersions. In Section 3, we give an explicit construction of an embedded doubling of an asymptotically conical shrinker. In Sections 4 and 5, we talk about the interior estimates and constructions of barrier functions to get the preservation and the improvement of $C^{2}$ bounds. In Section 6, we use the Ważewski box argument to finish the proof of Theorem 1 and conclude with the proof of Corollary 1.1.
@@ Acknowledgement
The authors are grateful to Toby Colding and Bill Minicozzi for their constant support and several inspiring conversations. They would like to thank Letian Chen, Otis Chodosh, Alec Payne, Max Stolarski, Ao Sun, Lu Wang, and Brian White for valuable discussions and comments, and to Alec Payne for bringing the work of Angenent, Ilmanen, and Velázquez to their attention and suggesting the application regarding fattening. During the project, Lee was partially supported by NSF Grant DMS 2005345 and Zhao was supported by NSF Grant DMS 1812142, NSF Grant DMS 1811267 and NSF Grant DMS 2104349.
@@ 2. Preliminaries facts and evolution equations
2.1. Analysis on asymptotically conical shrinkers
@@
There exists an $L^{2}_{W}$-orthonormal basis $\{h_{i}\}_{i\in\mathbb{N}}$ of $L^{2}_{W}(\Sigma)$ such that for all $i\in\mathbb{N},$ $h_{i}$ is an eigenfunction of $L$ with eigenvalue $\lambda_{i},$ that is, $Lh_{i}=-\lambda_{i}h_{i}.$111 We use the different sign convention about eigenvalues from the one used in [Sto]. The eigenvalues $\lambda_{i}$’s satisfy $\lambda_{1}<\lambda_{2}\leq\lambda_{3}\leq\cdots$ tending to +,$\infty$,+ ∞ , and each eigenspace is finite dimensional.
@@
$|h(x)|\leq C|x|^{2\left(\lambda+\delta+\frac{1}{2}\right)}.$
@@
Finally, we mention some weighted spaces that we will use following [CS]. Let $\tilde{r}$ be a fixed smooth function such that $\tilde{r}\geq 1$ and $\tilde{r}=|x|$ for $|x|\geq 2$. We consider the weighted norms
$\|f\|_{0;-\gamma}^{\rm hom}$ $:=\sup_{x\in\Sigma}\tilde{r}(x)^{\gamma}|f(x)|,\text{and}$
$[f]_{\alpha;-\gamma-\alpha}^{\rm hom}$ $:=\sup_{x,y\in\Sigma}\frac{1}{\tilde{r}(x)^{-\gamma-\alpha}+\tilde{r}(y)^{-\gamma-\alpha}}\frac{|f(x)-f(y)|}{|x-y|^{\alpha}}.$
The space $C^{0,\alpha}_{{\rm hom},-\gamma}$ is defined to collect all $f$ such that
$\|f\|_{C^{0,\alpha}_{{\rm hom},-\gamma}}:=\|f\|_{0;-\gamma}^{\rm hom}+[f]^{\rm hom}_{\alpha;-\gamma-\alpha}<\infty,$
and the space $C^{1,\alpha}_{{\rm hom},-\gamma}$ is defined to collect all $f$ such that
$\|f\|_{C^{1,\alpha}_{{\rm hom},-\gamma}}:=\sum_{j=0}^{1}\|(\nabla_{\Sigma})^{j}f\|_{C^{0,\alpha}_{{\rm hom},-j-\gamma}}<\infty.$
@@ 2.2. Evolution equation of immersions
$(\partial_{t}F_{0})(\varphi_{t}(x),t)+dF_{0}(\partial_{t}\varphi_{t}(x))(\varphi_{t}(x),t)=\sqrt{1-t}\,\partial_{t}{h}(x,t)\textbf{n}_{F}(x)-\frac{1}{2\sqrt{1-t}}\left(F(x)+{h}(x,t)\textbf{n}_{F}(x)\right).$
@@
$(\partial_{t}F_{0})(\varphi_{t}(x),t)+dF_{0}(\partial_{t}\varphi_{t}(x))(\varphi_{t}(x),t)+\frac{1}{2\sqrt{1-t}}F(x)=\left(\sqrt{1-t}\,\partial_{t}{h}(x,t)-\frac{1}{2\sqrt{1-t}}{h}(x,t)\right)\textbf{n}_{F}(x).$
As we assume that locally $\frac{1}{\sqrt{1-t}}F_{1}$ can be written as a graph over $F$, we have $\langle\textbf{n}_{F_{0}}(\varphi_{t}(x)),\textbf{n}_{F}(x)\rangle\neq 0.$ We write
@@
$|dF_{0}(\partial_{t}\varphi_{t}(x))|\leq\frac{\left|(\partial_{t}F_{0})(\varphi_{t}(x),t)+\frac{1}{2\sqrt{1-t}}F(x)\right|}{|\langle\textbf{n}_{F}(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle|}.$
@@
Setting $e^{\tau}=\frac{1}{1-t},$ we get $\partial_{t}=\frac{\partial\tau}{\partial t}\partial_{\tau}=\frac{1}{1-t}\partial_{\tau},$ and thus,
@@
$\hat{h}_{ab}$ $=\langle\partial_{a}F,\partial_{b}(h\,\textbf{n}_{F})\rangle+\langle\partial_{a}(h\,\textbf{n}_{F}),\partial_{b}F\rangle+\langle\partial_{a}(h\,\textbf{n}_{F}),\partial_{b}(h\,\textbf{n}_{F})\rangle=2hA_{ab}+\langle\partial_{a}h,\partial_{b}h\rangle+h^{2}A_{ia}A_{bj}g^{ij},\text{and}$
@@
$\|C_{1}(F,\nabla F,\nabla^{2}F,h,\nabla h)\|_{C^{0}}+\sum_{a,b}\|C_{2,ab}(F,\nabla F,\nabla^{2}F,h,\nabla h)\|_{C^{0}}+\|C_{3,ab}(F,\nabla F,\nabla^{2}F,h,\nabla h)\|_{C^{0}}\leq C_{0}\|h\|_{C^{1,\alpha}_{\rm hom,-1}}.$
Also, note that $\Delta^{g}\textbf{n}_{F}=-|A|^{2}\textbf{n}_{F}+\nabla^{g}H$ and $A_{ij,k}=A_{ik,j}$ in $\mathbb{R}^{n+1}$. Here, we use the notation $\textbf{H}=\Delta^{g}F$ and $H=g^{ij}A_{ij}=-\left\langle\textbf{H},\textbf{n}_{F}\right\rangle,$ so in normal coordinates, we have
@@ 3. Constructions of doubling of asymptotically conical shrinkers
As in (2.1), denoting the asymptotic cone of $\Sigma$ by $\mathcal{C}$, from [W14, CS], we know that outside a large compact set $\Sigma\cap\overline{B}_{R}$, we can parameterize $\Sigma\setminus B_{R}$ by
@@
over $\mathcal{C}\setminus B_{R}$, where $\theta$ is the parametrization for the link of $\mathcal{C}$ and n is the normal vector of $\mathcal{C}$ at $(r,\theta)$. We will modify $\Sigma$ in this region, and glue it with the compact part.
@@
$F_{3}(r,\theta)=\begin{cases}r\theta+\eta(\frac{r-R}{R})f(r,\theta)\textbf{n}&\text{if }R\leq r\leq 2R\\ r\cdot\frac{\left(\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)+\frac{1}{3}\right)\theta+\left(\frac{2}{3}-\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)\right)\mathbf{v}}{\left|\left(\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)+\frac{1}{3}\right)\theta+\left(\frac{2}{3}-\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)\right)\mathbf{v}\right|}&\text{if }r\geq 2R\end{cases}.$
@@
We can now glue $F$ with the original shrinker along $F(R,\cdot)$ with $\Sigma\cap\overline{B}_{R}$. As $\eta$ is a cutoff function, we have that the glued hypersurface is also smooth. From the constructions, the $F_{3}$ part lies in a hemisphere and controls the distance to origin. Thus, we know that $F$ is globally an embedding, and after reflecting it with respect to the hyperplane $\{x\in\mathbb{R}^{n+1}\,|\,\langle x,\mathbf{v}\rangle=B+R\},$ we get a closed embedded hypersurface.
@@ 4. Interior estimates
$\|h\|_{C^{0,\alpha}_{\rm hom,-1}(\Sigma\times[0,\tau_{1}])}$ $<\varepsilon,$
$\|h(\tau)\|_{L^{2}_{W}(\Sigma)}$ $\leq\mu e^{-\lambda_{*}(\tau+\tau_{0})}\,\,\text{ for all}\,\,\tau\in[0,\tau_{1}],\text{ and}$
$\sum_{l=0}^{4n+3}\left|\nabla^{l}h\right|_{g}(x,0)$ $\leq C_{0}\bar{p}e^{-\lambda_{*}\tau_{0}}\,\,\text{ for all}\,\,x\in\left\{\frac{|F|^{2}}{4}<2\Gamma e^{\tau_{0}}\right\}\text{ and }\tau=0$
for some $\lambda_{*}>0.$ Then for $\Gamma>\Gamma_{0}(n,\Sigma,F)$, $\varepsilon<\varepsilon_{0}(n,\Sigma,F,\Gamma),$ and $\tau_{0}>C(n,\Sigma,F,\lambda_{*},\Gamma,C_{0}\bar{p})$, there exists $C_{1}=C_{1}(n,\Sigma,F,\lambda_{*},\Gamma)<\infty$ such that
$\sum_{l=0}^{2}|\nabla^{g}h|_{g}(x,\tau)\leq C_{1}(\mu+C_{0}\bar{p})e^{-\lambda_{*}(\tau+\tau_{0})}\,\,\text{ for all}\,\,(x,\tau)\in\left\{\frac{|F|^{2}}{4}<\Gamma e^{\tau_{0}}\right\}\times[0,\tau_{1}].$
@@
This helps us to get rid of the unbounded coefficient term $-\frac{1}{2}\langle F,\nabla^{g}f\rangle$ in $L$. Also, note that the relation $0\leq\left|\nabla^{g}(\frac{|F|^{2}}{4})\right|^{2}\leq\frac{|F|^{2}}{4}$ implies that for $s_{1}<s_{2},$
@@
Moreover, from [CS, Lemma 3.8], if we denote the link of the asymptotic cone of $\Sigma$ by $\mathcal{C}$, by parameterizing the shrinker outside a compact set by $F:(R,\infty)\times\Gamma\to\Sigma$ and denoting the family of diffeomorphisms generated by $X_{t}=\frac{1}{2(-t)}x^{T}$ by $\Phi_{t}$, we can define $\tilde{\Phi}_{t}=F^{-1}\circ\Phi_{t}\circ F$ and $\phi_{t}:(R,\infty)\times\Gamma\to(R,\infty)\times\Gamma,\,(r,\omega)\to((-t)^{-\frac{1}{2}}r,\omega)$ with the following estimates.
[…]
$d_{g_{\mathcal{C}}}(\tilde{\Phi_{t}}(t,\theta),\phi_{t}(r,\theta))\lesssim\frac{1}{\sqrt{-t}\cdot r},$
$|D^{(j)}(\tilde{\Phi}-\phi_{t})|(r,\theta)\lesssim\frac{1}{\sqrt{-t}\cdot r^{1+j-\eta}},$
for $j\geq 1$ and $\eta>0$.
As in [CS], we recall a non-standard interior Schauder estimates from [Bra69].
[…]
Suppose $B_{2}\in\mathbb{R}^{n}$ and there are $C^{0,\alpha}$ functions $a_{ij},b_{i},c,u,f\colon B_{2}\times[-2,0]\to\mathbb{R}$ such that
$\frac{\partial}{\partial t}u=a_{ij}D_{ij}^{2}u+b_{i}D_{i}u+cu+f.$
If the system is uniformly parabolic in the sense that $a_{ij}\xi_{i}\xi_{j}\geq\lambda|\xi|^{2}$ for some $\lambda>0$ and
$\sup_{t\in[-2,0]}\left(\|a_{ij}(\cdot,t)\|_{C^{0,\alpha}(B_{1})}+\|b_{i}(\cdot,t)\|_{C^{0,\alpha}(B_{1})}+\|c(\cdot,t)\|_{C^{0,\alpha}(B_{1})}\right)\leq\Lambda$
for some $\Lambda>0,$ then for $T\in(-1,0],$
$\sup_{t\in[-1,T]}\|u(\cdot,t)\|_{C^{2,\alpha}(B_{1})}\leq C\sup_{t\in[-2,T]}\left(\|u(\cdot,t)\|_{C^{0,\alpha}(B_{2})}+\|f(\cdot,t)\|_{C^{0,\alpha}(B_{2})}\right)$
for some $C=C(n,\alpha,\lambda,\Lambda)$.
Combining Lemma 4, Lemma 4 and Proposition 4 implies the large scale interior estimates.
@@
$|h|_{g}$ $\leq C_{0}e^{-k(\tau+\tau_{0})}\left(\frac{|F|^{2}}{4}(x)\right)^{k+\frac{1}{2}}\,\,\text{on}\,\,\left\{(x,\tau)\in\Sigma\times(0,\tau_{1}):\,\Gamma e^{\tau_{0}}<\frac{|F|^{2}}{4}<\gamma e^{\tau+\tau_{0}}\right\},\text{ and}$
$|\nabla^{l}h|_{g}(x,\tau_{0})$ $\leq C_{0}e^{-k\tau_{0}}\left(\frac{|F|^{2}}{4}(x)\right)^{k+\frac{1}{2}-\frac{l}{2}}\,\,\text{for all}\,\,l\in\{0,1,2,3\},\,\,x\in\left\{\Gamma e^{\tau_{0}}<\frac{|F|^{2}}{4}<\gamma e^{\tau+\tau_{0}}\right\}.$
@@
$|\nabla^{l}h|_{g}(x,\tau)\leq C_{1}C_{0}e^{-k(\tau+\tau_{0})}\left(\frac{|F|^{2}}{4}(x)\right)^{k+\frac{1}{2}-\frac{l}{2}}\,\,\text{on}\,\,\left\{(x,\tau)\in\Sigma\times[0,\tau_{1}]:\,4\Gamma e^{\tau_{0}}\leq\frac{|F|^{2}}{4}(x)\leq\frac{1}{4}\gamma e^{\tau+\tau_{0}}\right\}$
for all $l\in\{0,1,2\}.$
Similarly, in the outer region, we have the following estimate.
@@
$|h|_{g}(x,\tau)$ $\leq\delta|F|\,\,\text{on}\,\,\left\{(x,t)\in\Sigma\times[0,\tau_{1}]:\,\gamma e^{\tau+\tau_{0}}<\frac{|F|^{2}}{4}(x)\right\},$
$|\nabla^{l}h|_{g}(x,0)$ $\leq\delta|F|^{1-l},\,\,l\in\{0,1,2,3\}\,\,\text{on}\,\,x\in\left\{\gamma e^{\tau_{0}}\leq\frac{|F|^{2}}{4}\right\},\text{ and}$
@@
$|\nabla^{l}h|_{g}(x,\tau)\leq C\delta|F|^{1-l},l\in\{0,1,2\}\,\,\text{on}\,\,\left\{(x,\tau)\in\Sigma\times[0,\tau_{1}]:\,4\gamma e^{\tau+\tau_{0}}<\frac{|F|^{2}}{4}(x)\right\}.$